OpenStudy (anonymous):

Find the equation of the circle. x^2+y^2+Dx+Ey+F=0 that passes through the points: (2,0)(-1,-3)(5,-3)

OpenStudy (anonymous):

simply input the points to the equation. when you put (2,0) you got equation1. : (2)^2 +2D +F = 0 => 2D+F=-4 Then the (-1,-3) and (5,-3) so that you get 2 more equations : equation 2: -D-3E+F=-10 and equation 3: 5D-3E+F =-34 then if you do elimination on equation 2 and 3, so that you'll get -6D=24 . Therefore D= -4 then put the value of D on equation 1. you'll get F=-4+8 = 4. then put the value of D and F to equation 2 or 3 and you'll get E=-18. Therefore the equation of the circle is : x^2 + y^2 - 4x - 18y +4 = 0

OpenStudy (anonymous):

i think i got the value of E wrong. the value of E should be 6 . then the right equation is ^2 + y^2 - 4x - 6y +4 = 0

OpenStudy (anonymous):

ups sorry ,typo. the equation is x^2 + y^2 - 4x + 6y +4 = 0.

OpenStudy (anonymous):

WOw this makes a lot more sense than my book! Thank you!

OpenStudy (anonymous):

you're welcome :)